Can youi help me with this problem? I am trying to simplify this rational expression:
x^3-8/4-x^2

Answer by rothauserc(4718)

(x^3 - 8) / (4 - x^2) = ((x - 2) * (x^2 +2x + 4)) / (-(x+2)*(x-2)) =
(x^2 +2x + 4) / (-(x+2))